A force of 100 N makes an angle of θ with the x-axis and has a scalar y-component of 30 N. Find both the scalar x-component of the force and the angle θ.
1
Expert's answer
2020-04-01T10:14:41-0400

"\sin\theta=\frac{30}{100}=0.3\to \theta=17.46°"


x-component of the force


"100\cdot\cos17.46°=95.4 N"